# Semantic Edge Computing and Semantic Communications in 6G Networks: A Unifying Survey and Research Challenges

### v2

@@ Abstract
Semantic Edge Computing (SEC) and Semantic Communications (SemComs) have been proposed as viable approaches to achieve real-time edge-enabled intelligence in sixth-generation (6G) wireless networks. On one hand, SemCom leverages the strength of Deep Neural Networks (DNNs) to encode and communicate the semantic information only, while making it robust to channel distortions by compensating for wireless effects. Ultimately, this leads to an improvement in the communication efficiency. On the other hand, SEC has leveraged distributed DNNs to divide the computation of a DNN across different devices based on their computational and networking constraints. Although significant progress has been made in both fields, the literature lacks a systematic view to connect both fields. In this work, we fulfill the current gap by unifying the SEC and SemCom fields. We summarize the research problems in these two fields and provide a comprehensive review of the state of the art with a focus on their technical strengths and challenges.
@@ 1 Introduction
For decades, wireless communication is achieved by encoding information in binaries, introducing redundancy in bitstreams for error correction, and transmitting bits through the unreliable physical layer [1]. However, such bit level communication paradigm is constrained by the Shannon capacity limit and cannot meet the growing need of data transmission throughput in six-generation (6G) wireless networks [2]. For example, for a video-streaming application, sending frames at 120 Hz frame rate with 8K resolution and 8 bits resolution per pixel requires more than 30 Gbps of bandwidth while the IEEE 802.11 ax standard can only provide up to 9.6 Gbps speed.
@@
To this end, the field of Semantic Communication (SemCom) has investigated data-driven Deep Learning (DL)-based techniques for transmitting signals at the “semantic” level. Specifically, SemCom trains a Deep Neural Network (DNN)-based source and channel encoder and decoder deployed separately at the transmitter and the receiver side, to extract and decode the semantic information that can be directly transmitted over the air [2]. Compared to bit-level communication system that separately optimize modulation, source and channel coding schemes, SemCom can achieve more efficient communication through transmission of compressed semantic information and learning-driven joint optimization of modulation and coding schemes.
@@
On the other hand, many emerging DL-based technologies, envisioned as potential applications of 6G networks – i.e., metaverse, smart city, autonomous driving, etc – require a real-time performance of DNNs which demand significant computational resources. Conventional edge computing that offload the entire computation to powerful edge servers can result in substantial latency for data transmission. To address this challenge, Semantic Edge Computing (SEC) is proposed to alleviate the computation and communication overhead of DL applications by distributing the workload [3]. The key idea is to split the computation of DNNs across multiple devices based on their computational and networking constraints and transmit latent representations – i.e., the “semantic” extracted by the DNN– through 6G networks to accelerate the inference. In stark opposition to conventional edge computing where the entire raw data is offloaded to edge, SEC reduce the communication overhead by leveraging a bottleneck structure at the splitting point of DNN to compress the latent representations, hence achieving lower latency [4, 5, 6].
@@
While these two fields are derived independently for different purposes, they share notable similarities in their goals and methodologies, which is further discussed in Section 2. For instance, both SEC and SemCom leverage DNNs distributed on separate devices for extracting and compressing semantic information that need to be transmitted in wireless networks. SEC primarily focuses on optimizing computation, while SemCom emphasizes the resilience of semantics in wireless channels. Merging techniques in both fields is promising to achieve more efficient and reliable DL applications in 6G networks.
@@
However, the key issue is that despite the striking similarities in goals between the two, the fields of SEC and SemCom have largely ignored each other. For example, although there are some recent work in SEC which considered adaptive latent feature compression to varying channel conditions [7], they leveraged conventional network protocols such as Wi-Fi to transfer latent representations and neglect advanced coding scheme in SemCom. Similarly, while prior work has leveraged Information Bottleneck (IB) theory to compress the encoded semantics [8, 9], most of the work in SemCom used DNNs with custom-tailored fixed architectures to build communication systems without considering recent feature compression techniques in SEC. To fill the existing gap, this paper provides an unified view of SEC and SemCom to demystify their similarities and summarize the latest breakthroughs in both areas. Our work is orthogonal to existing literature reviews [3, 2] since the unified overview helps both research fields to discover the similarity and adopt new perspective from the other community.
@@
Paper Organization. Section 2 introduces the background and an overview and discusses the relation between SEC and SemCom. Section 3 and Section 4 summarize the recent progress in SEC and SemCom, respectively. We conclude our paper by discussing the open challenges and future directions in Section 5 and providing the summary in Section 6.
2 Overview of Semantic Edge Computing and Semantic Communication
@@ 2.1 Semantic Edge Computing
6G networks are envisioned to facilitate many DL-driven technologies such as autonomous driving and Internet of Things (IoT) [10], resulting in an increasing need to deploy data-driven algorithms based on DL on resource-constrained devices such as mobile phones. Mobile DNN, aiming at creating hardware-efficient architectures, often lead to a notable performance degradation [11]. Edge computing preserves the DNN performance by offloading the entire computation to powerful edge servers [12], but can result in a large latency due to the excessive data transmission.
@@
SEC, also referred to as “split computing” [6] or “DNN partitioning” [13] in prior art, has been proposed to accelerate DNN inference by leveraging computational resources across devices in a cooperative manner. The key idea is to send only the compressed representation of the information – i.e., the “semantic” – rather than the input to minimize the latency. As depicted in Figure 1, SEC divides a DNN into two parts, which are executed on mobile and edge devices separately. The light-weight DNN executed by the mobile device processes the raw input and generates latent representations which contain the semantics of the input. On the other hand, a larger DNN running on the edge server will take the semantics and generate the output of the task inference. The data is transferred between the mobile and edge through wireless communication. Compared to mobile DNNs, SEC can leverage the computational power on edge device to achieve a better DL performance. SEC also achieves less latency compared to conventional edge computing by transmitting semantics that have smaller size than raw inputs [3].
Formally, a DNN $\mathcal{M}$ can be modeled as a Markov chain $\mathcal{M}:X\mapsto L_{1}\mapsto L_{2}\mapsto\cdots\mapsto L_{i}\mapsto\cdots\mapsto Y$ where $X$, $Y$ and $L_{i}$ are input data, output data, and the latent representation of $i$-th hidden layer. We assume the Markov chain is split into a head model $\mathcal{H}:X\mapsto\cdots\mapsto L_{i}$ and a tail model $\mathcal{T}:L_{i}\mapsto\cdots\mapsto Y$, deployed separately on the mobile and edge device. However, simply splitting the model at the $i$-th layer may result in excessive communication burden as the latent representation $L_{i}$ can have large data size without compression. For example, the intermediate representation in early layers of ResNet [14] can go up to 5x larger than the original input data. On the other hand, compressing latent features can lead to significant performance loss. Therefore, a fundamental research question in SEC is to achieve optimal feature compression without compromising DNN performance [6, 15, 16]. In addition, splitting DNNs at different layers can result in different computation-latency trade-off [17]. Therefore, optimizing the splitting point of DNN based on different computational and networking constraints is a key factor to SEC [18]. In summary, SEC aims to achieve the following goals: (1) optimize the computation load across mobile and edge devices; and (2) compress the semantic information for inference that needs to be transmitted to reduce the communication latency.
@@ 2.2 Semantic Communication
The first level – corresponding to the classical communication systems and called the bit level – is designed to reliably transfer the symbolic representation of data from the transmitter to the receiver. There is a vast research investigating this type of communication which usually use conventional independent and separately-optimized blocks to develop the system [19, 20, 21]. In classical communication systems, a transmitter, which aims to send a message $s$, encodes it into $x$ using separate source and channel coding blocks and transmits it via a channel. The intended receiver receives a distorted signal $y$, and tries to accurately decode the message $\hat{s}$ in terms of the bit-level metrics. Assuming the channel matrix to be $H$, the received signal can be derived as $y=Hx+\epsilon$, where $\epsilon$ denotes the channel noise.
@@
The second type of systems – called the semantic-level systems – are focused on extracting the semantics of data and often use DNNs both at the transmitter and receiver which are jointly trained in an end-to-end manner incorporating the channel model [22, 23, 24, 25]. Such systems can break the Shannon’s limit by transmitting only the semantic meaning behind the message instead of raw bits. As such, as long as the receiver can understand and infer the expected important information correctly, the loss of semantically-irrelevant and redundant content can be tolerated. Note that while semantic and channel coding blocks are shown separately in Figure 2, both functionalities can be merged together which is also referred to as Joint Source-Channel Coding (JSCC) in the literature. More specifically, the semantic encoder extracts semantic features using the integrated Knowledge Base (KB) from the input $s$ and filters out the redundant information. Then, the channel encoder generates the channel symbols $x$, which are sent through the channel. During transmission, the symbols $x$ are distorted according to the semantic noise accounting for the physical noise, semantic mismatch between the source and destination KBs, ambiguity and interpretation errors. Semantic noise extends the concept of physical noise in classical communication to any noise causing semantic distortion. At the receiver, the channel decoder estimates the semantically-corrupted features using the received symbols $y$, and the semantic decoder recovers the semantically-similar message $\hat{s}$ based on the destination KB by maximizing the semantic similarity metrics [26]. Therefore, both the transmitter and receiver require to share and continuously update a KB using the knowledge from the environment.
@@
Most of the existing work view the entire SemCom system as a DL-based reconstruction system, and use an encoder-decoder architecture for the DNN with the bottleneck representing the encoded data to be transmitted over the channel. Also, a stochastic model representing the channel effects is incorporated between the encoder and decoder DNN, and the communication system is trained in an end-to-end manner to maximize the similarity metrics between the encoder input and the decoder output. As a result, the DNNs learn the channel distortion patterns and make the encoded symbols robust to wireless impairments. In general, SemCom focuses on: (1) how to effectively extract the semantic information and encode such semantics into channel symbols; and (2) how to make the semantics robust to the noisy wireless channel.

### v3

@@ Abstract
Semantic Edge Computing (SEC) and Semantic Communications (SemComs) have been proposed as viable approaches to achieve real-time edge-enabled intelligence in sixth-generation (6G) wireless networks. On one hand, SemCom leverages the strength of Deep Neural Networks (DNNs) to encode and communicate the semantic information only, while making it robust to channel distortions by compensating for wireless effects. Ultimately, this leads to an improvement in the communication efficiency. On the other hand, SEC has leveraged distributed DNNs to divide the computation of a DNN across different devices based on their computational and networking constraints. Although significant progress has been made in both fields, the literature lacks a systematic view to connect both fields. In this work, we fill the current gap by unifying the SEC and SemCom fields. We summarize the research problems in these two fields and provide a comprehensive review of the state of the art with a focus on their technical strengths and challenges.
@@ 1 Introduction
For decades, wireless communication has been achieved by encoding information in binaries, introducing redundancy in bitstreams for error correction, and transmitting bits through the unreliable physical layer [1]. However, such a bit-level communication paradigm is constrained by the Shannon capacity limit and cannot meet the growing need of data transmission throughput in sixth-generation (6G) wireless networks [2]. For example, for a video-streaming application, sending frames at 120 Hz frame rate with 8K resolution and 8 bits resolution per pixel requires more than 30 Gbps of bandwidth while the IEEE 802.11 ax standard can only provide up to 9.6 Gbps speed.
@@
To this end, the field of Semantic Communication (SemCom) has investigated data-driven Deep Learning (DL)-based techniques for transmitting signals at the “semantic” level. Specifically, SemCom trains a Deep Neural Network (DNN)-based source and channel encoder and decoder deployed separately at the transmitter and receiver sides, to extract and decode the semantic information that can be directly transmitted over the air [2]. Compared to bit-level communication systems that separately optimize modulation, source and channel coding schemes, SemCom can achieve more efficient communication through transmission of compressed semantic information and learning-driven joint optimization of modulation and coding schemes.
@@
On the other hand, many emerging DL-based technologies, envisioned as potential applications of 6G networks – i.e., metaverse, smart city, autonomous driving, etc – require a real-time performance of DNNs which demand significant computational resources. Conventional edge computing that offloads the entire computation to powerful edge servers can result in substantial latency for data transmission. To address this challenge, Semantic Edge Computing (SEC) is proposed to alleviate the computation and communication overhead of DL applications by distributing the workload [3]. The key idea is to split the computation of DNNs across multiple devices based on their computational and networking constraints and transmit latent representations – i.e., the “semantic” extracted by the DNN– through 6G networks to accelerate the inference. In stark opposition to conventional edge computing where the entire raw data is offloaded to the edge, SEC reduces the communication overhead by leveraging a bottleneck structure at the splitting point of DNN to compress the latent representations, hence achieving lower latency [4, 5, 6].
@@
While these two fields are derived independently for different purposes, they share notable similarities in their goals and methodologies, which are further discussed in Section 2. For instance, both SEC and SemCom leverage DNNs distributed on separate devices for extracting and compressing semantic information that needs to be transmitted in wireless networks. SEC primarily focuses on optimizing computation, while SemCom emphasizes the resilience of semantics in wireless channels. Merging techniques from both fields is promising to achieve more efficient and reliable DL applications in 6G networks.
@@
However, the key issue is that despite the striking similarities in goals between the two, the fields of SEC and SemCom have largely ignored each other. For example, although there is some recent work in SEC which considered adaptive latent feature compression to varying channel conditions [7], it leveraged conventional network protocols such as Wi-Fi to transfer latent representations and neglected advanced coding scheme in SemCom. Similarly, while prior work has leveraged Information Bottleneck (IB) theory to compress the encoded semantics [8, 9], most work in SemCom used DNNs with custom-tailored fixed architectures to build communication systems without considering recent feature compression techniques in SEC. To fill the existing gap, this paper provides a unified view of SEC and SemCom to demystify their similarities and summarize the latest breakthroughs in both areas. Our work is orthogonal to existing literature reviews [3, 2] since the unified overview helps both research fields to discover the similarity and adopt a new perspective from the other community.
Scope and Contributions. In this survey, we focus on the technologies that aim to achieve edge intelligence through collaborative execution of DNNs during inference time. Our scope deliberately excludes federated learning and split learning which focus on distributing DNN computation during the training phase.
Several existing studies address related topics such as split computing, collaborative inference, and DNN partitioning [3, 10, 11, 12] for distributing computation during inference to accelerate DL applications. For example, Matsubara et al. [3] summarized the split computing methods to accelerate DNN applications in edge computing. Ren et al. [10] focused on optimizing the DNN computation allocation in the device-edge-cloud collaborative environment. Wang et al. [11] reviewed the security and privacy challenges of this distributed DNN system during the communication across devices. Zhang et al. [12] summarized the training, inference and knowledge transfer technologies in mobile-edge-cloud computing.
Despite overlapping scopes, these topics typically emphasize different aspects of edge computing and lack a comprehensive perspective. For example, split computing primarily focuses on optimizing DNN architecture and training methods to minimize computation on resource-limited devices while preserving task performance. In contrast, cooperative inference or DNN partitioning concentrates on resource allocation within multi-device collaborative systems for deploying pre-trained, off-the-shelf DNNs. Compared to other work that solely focuses on split computing [3] or DNN partitioning [10, 11], our survey aims to provide a more holistic view of these interconnected approaches. As such, we introduce the term “Semantic Edge Computing (SEC)” to encompass these concepts, acknowledging their shared foundation of transmitting latent representations (i.e., “semantics”) extracted by DNNs for task execution.
In addition, SemCom is recognized as a promising approach to enable DL deployment at the edge. Several surveys explore different aspects of SemCom. For instance, Yang et al. [13] reviewed the theory development and proposed a categorization of SemCom technologies. Luo et al. [2] explained the SemCom concept and studied the DL progress in this field. Guo et al. [14] highlighted the security and privacy challenges of SemCom in a multi-agent setting while Getu et al. [15] reviewed the current progress with an emphasis on 6G applications. However, existing surveys overlook related research in SEC, which could potentially enhance SemCom through the integration of sophisticated feature extraction techniques and system optimization strategies. Our work separates itself by providing a new framework that connects SemCom to SEC with a detailed discussion on the methodology and challenges. Table 1 summarizes our main contribution and distinction from other relevant surveys.
@@
Paper Organization. Section 2 introduces the background and an overview to discuss the relation between SEC and SemCom. Section 3 and Section 4 summarize the recent progress in SEC and SemCom, respectively. We conclude our paper by discussing the limitations and future directions in Section 5 and providing a summary in Section 6.
Table 1: A Comparison between Relevant Survey and Our Work
Category Year Key focus and contributions Our contributions and distinctions
SemCom 2022 [13] It provides a roadmap of SemCom theory and categorization of SemCom technologies with a discussion on their benefit and limitations. $\bullet$ We categorize SemCom, and propose a new framework to connect SemCom and SEC. $\bullet$ We propose designing methods to this framework, addressing its advantages and challenges. $\bullet$ We provide a comprehensive literature review of recent progress in SemCom technologies.
2022 [2] It explains the SemCom concept and reviewed recent deep learning in this emerging field.
2025 [15] It summarizes the SemCom framework and current research with an emphasis on 6G applications.
2024 [14] It introduces a three-layer framework for multi-agent SemCom and explores security and privacy issues.
SEC 2022 [3] It reviews tasks, technologies and models of split computing and early exiting in device-edge computing $\bullet$ We summarize SEC, including split computing, cooperative inference, and DNN partitioning. $\bullet$ We categorize SEC based on their technical focus, revealing the connection to SemCom. $\bullet$ We provide a comprehensive review of the methodologies, technical challenges of SEC. $\bullet$ We discuss limitations of current SemCom and SEC and potential future research directions.
2023 [10] It summarizes device-device, device-edge, edge-cloud collaborative inference with a focus on optimization technologies
2024 [11] It reviews the security and privacy challenges and technologies in edge- cloud DNN partitioning
2025 [12] It focuses on model training, inference and updating in mobile-edge-cloud computing, with an emphasis on compression, partitioning and knowledge transfer technologies
2 Background of Semantic Edge Computing and Semantic Communication
2.1 Edge Intelligence in 6G Network
6G Network for Intelligence. The large network capacity and massive connectivity of 6G networks is expected to facilitate many DL-driven applications such as Integrated Sensing and Communication (ISAC) [16, 17, 18], metaverse [19, 20] and autonomous driving [21, 22, 23]. ISAC can achieve object tracking, localization, and activity recognition using the Channel State Information (CSI) [16] or Beamforming Feedback Information (BFI) [18] of the signal which carries data traffic. Compared to conventional sensing methods that need independent infrastructure, ISAC leverages the existing communication signals and spectrum to perform sensing, creating a unified system where communication and sensing capabilities are intrinsically merged within the same hardware and spectrum resources. In ISAC, DNNs are essential for processing and interpreting the complex patterns within CSI or BFI to distinguish between different objects, movements, and environmental changes. The large bandwidth and high-volume data transmission capabilities of 6G networks can enhance the available object information from carrier signals, enabling significantly improved sensing resolution.
In addition, metaverse relies heavily on DL algorithms for a continuous 3D environment rendering, human pose estimation and spatial tracking while autonomous driving leverages DNNs for object detection, dynamic path planning, and decision-making based on massive sensor data streams. All these applications require a massive data transmission and processing. As such, the network capacity of 6G networks can enable seamless real-time interactions in immersive metaverse environments while providing the instantaneous data processing and communication required for safe, coordinated autonomous vehicle networks.
Intelligence for 6G Network. On the other hand, 6G is conceived as an intelligent network where DL is integrated as a foundational layer, enabling adaptive, self-optimized and massive communication. For instance, Radio Frequency Machine Learning System (RFMLS) is a new signal processing paradigm that applies modern machine learning techniques and DNNs to the Radio Frequency (RF) spectrum domain. It leverages the strength of DNN to process RF waveforms, identify different signals and detect anomalies [24, 25, 26]. As such, it enables efficient spectrum management where the transmitter can dynamically adjust parameters such as power levels, modulation schemes, and frequency allocations based on real-time spectrum conditions and varying traffic demands, significantly improving spectral efficiency and network performance.
In addition, Open Radio Access Network (Open-RAN) [27, 28, 29], which aims to transform traditional RAN into open, standardized, and interoperable elements, is expected to be the foundation of 6G network. Compared to conventional RANs that are highly integrated and vendor-specific, Open-RAN creates independent, modularized interfaces for flexible connections to different network functions. It uses RAN Intelligent Controllers (RICs) to seamlessly integrate various DL solutions for autonomous network resource management and performance optimization.
Edge Intelligence. However, achieving intelligence in 6G networks faces critical computation and communication challenges. Research has demonstrated that DNNs harvest performance from scaling-up model size [30] while many 6G technologies need to be deployed on resource-constrained devices. A possible solution is to create lightweight, hardware-efficient DNN architectures. However, these compressed models typically exhibit substantial performance degradation. For example, MobileNetV2 suffers more than 3% accuracy loss on ImageNet classification compared to ResNet50 [31]. While offloading computation to powerful cloud servers represents a viable solution to overcome local resource constraints, this approach introduces massive communication overhead. For instance, metaverse requires the DL algorithm to continuously process the high quality video streams to create an immersive AR/VR environment. Offloading large volumes of data to edge/cloud can result in significant latency. To this end, SEC and SemCom are proposed to bring intelligence to the edge by distributing computation across multiple devices and minimize the inter-device communication overhead.
@@ 2.2 Semantic Edge Computing
To address the computation and communication challenge in deploying DNNs at edge, prior art proposed “DNN partitioning” [32] and “split computing” [6] to accelerate DNN inference by leveraging computational resources across devices in a cooperative manner. The key idea of DNN partitioning is to divide DNN layers across multiple devices to overcome the computation burden that no single resource-constrained device could handle. Compared to conventional edge offloading that transmits raw input, the collaborative inference requires the system to send latent representations – i.e., the “semantic” – generated at the DNN splitting point. DNN partitioning concentrates on complex optimization with various conditions such as network bandwidth, device computational capabilities, energy consumption, and latency requirements, with the goal of minimizing end-to-end inference time and ensuring efficient resource utilization.
However, simply splitting DNNs may result in excessive communication burden as the latent representation can have large data size without compression. For example, the intermediate representation in early layers of ResNet can go up to 5x larger than the original input data [33]. As such, research in split computing focuses on DNN architecture design and advanced training mechanism to compress the latent representation at the splitting point while maintaining sufficient semantic information for task performance.
Despite their high relevance, existing literature in split computing and DNN partitioning focus on narrow perspectives. In this survey, we use “Semantic Edge Computing (SEC)” to unify these overlapping topics and summarize the state of the art in each field. We adopt the term “semantic” to distinguish this approach from traditional edge computing paradigms that offload entire computational tasks to remote servers and transmit raw input data for processing. In contrast, SEC selectively offloads specific computational components while transmitting semantically meaningful features extracted by intermediate layers, rather than unprocessed data. Formally, SEC refers to a designing pattern that a DNN $\mathcal{M}$ can be modeled as a Markov chain $\mathcal{M}:X\mapsto L_{1}\mapsto L_{2}\mapsto\cdots\mapsto L_{i}\mapsto\cdots\mapsto Y$. $X$, $Y$ and $L_{i}$ are input data, output data, and the latent representation of $i$-th hidden layer. Without loss of generality we assume the Markov chain is split into a head model $\mathcal{H}:X\mapsto\cdots\mapsto L_{i}$ and a tail model $\mathcal{T}:L_{i}\mapsto\cdots\mapsto Y$, deployed separately on the mobile and edge device.
As summarized in Figure 1, compared to mobile DNNs, SEC can leverage the computational power on edge devices to achieve better DL performance. SEC also achieves less latency compared to conventional edge computing by transmitting semantics that have a smaller size than raw inputs [3]. SEC aims to achieve the following goals: (1) optimize the computation load across mobile and edge devices; and (2) compress the semantic information for inference that needs to be transmitted to reduce the communication latency.
@@ 2.3 Semantic Communication
The first level – corresponding to the classical communication systems and called the bit level – is designed to reliably transfer the symbolic representation of data from the transmitter to the receiver. There is vast research investigating this type of communication which usually uses conventional independent and separately-optimized blocks to develop the system [34, 35, 36]. In classical communication systems, a transmitter, which aims to send a message $s$, encodes it into $x$ using separate source and channel coding blocks and transmits it via a channel. The intended receiver receives a distorted signal $y$, and tries to accurately decode the message $\hat{s}$ in terms of the bit-level metrics. Assuming the channel matrix to be $H$, the received signal can be derived as $y=Hx+\epsilon$, where $\epsilon$ denotes the channel noise.
@@
The second type of systems – called the semantic-level systems – are focused on extracting the semantics of data and often use DNNs both at the transmitter and receiver which are jointly trained in an end-to-end manner incorporating the channel model [37, 38, 39, 40]. Such systems can break the Shannon’s limit by transmitting only the semantic meaning behind the message instead of raw bits. As such, as long as the receiver can understand and infer the expected important information correctly, the loss of semantically irrelevant and redundant content can be tolerated. Note that while semantic and channel coding blocks are shown separately in Figure 2, both functionalities can be merged together which is also referred to as Joint Source-Channel Coding (JSCC) in the literature. More specifically, the semantic encoder extracts semantic features using the integrated Knowledge Base (KB) from the input $s$ and filters out the redundant information. Then, the channel encoder generates the channel symbols $x$, which are sent through the channel. During transmission, the symbols $x$ are distorted according to the semantic noise accounting for the physical noise, semantic mismatch between the source and destination KBs, ambiguity and interpretation errors. Semantic noise extends the concept of physical noise in classical communication to any noise causing semantic distortion. At the receiver, the channel decoder estimates the semantically corrupted features using the received symbols $y$, and the semantic decoder recovers the semantically similar message $\hat{s}$ based on the destination KB by maximizing the semantic similarity metrics [41]. Therefore, both the transmitter and receiver require to share and continuously update a KB using the knowledge from the environment.
@@
Most of the existing work views the entire SemCom system as a DL-based reconstruction system, and use an encoder-decoder architecture for the DNN with the bottleneck representing the encoded data to be transmitted over the channel. Also, a stochastic model representing the channel effects is incorporated between the encoder and decoder DNN, and the communication system is trained in an end-to-end manner to maximize the similarity metrics between the encoder input and the decoder output. As a result, the DNNs learn the channel distortion patterns and make the encoded symbols robust to wireless impairments. In general, SemCom focuses on: (1) how to effectively extract the semantic information and encode such semantics into channel symbols; and (2) how to make the semantics robust to the noisy wireless channel.